# Machine learning for advancing low-temperature plasma modeling and simulation

### v1

@@ Abstract
Machine learning has had an enormous impact in many scientific disciplines. Also in the field of low-temperature plasma modeling and simulation it has attracted significant interest within the past years. Whereas its application should be carefully assessed in general, many aspects of plasma modeling and simulation have benefited substantially from recent developments within the field of machine learning and data-driven modeling. In this survey, we approach two main objectives: (a) We review the state-of-the-art focusing on approaches to low-temperature plasma modeling and simulation. By dividing our survey into plasma physics, plasma chemistry, plasma-surface interactions, and plasma process control, we aim to extensively discuss relevant examples from literature. (b) We provide a perspective of potential advances to plasma science and technology. We specifically elaborate on advances possibly enabled by adaptation from other scientific disciplines. We argue that not only the known unknowns, but also unknown unknowns may be discovered due to an inherent propensity to spotlight hidden patterns in data.
@@ 1 Introduction
Low-temperature plasmas (LTPs) consist of a mixture of neutral and charged species in thermal non-equilibrium. The governing discharge physics and chemistry lead to a zoo of (excited) species interacting with bounding surfaces. Their modification (e.g., activation, etching, functionalization) is exploited in the frame of plasma processing (e.g., plasma-enhanced etching, deposition, catalysis). LTPs also render a dominant part of the many steps required in semiconductor device manufacturing. This may involve a diverse portfolio of plasma discharges, chemical precursors, surface materials, and process recipe steps to achieve desired deposit/etch patterns.[1, 2] It may as well regard indirect plasma discharges for example in the generation of extreme ultraviolet light for state of the art 13.5 nm wavelength lithography, required for device feature scales down to a few 10 nm and beyond (for instance termed 3 nm nodes).[3] Modeling and numerical simulation contribute an integral component in the design and optimization of related plasma processes. More recently in the past decade, data-driven methods such as machine learning (ML), deep learning (DL), and artificial neural networks (ANNs), among other computational statistics methods – which may be subsumed under the umbrella term artificial intelligence (AI) – have also (re-)gained enormous interest. Although these fields have seen a renaissance and have continued to rapidly develop in the past decade, many fundamental concepts have been well established in computational statistics and related disciplines. The past and present interest in the context of plasma processing may be attributed to the potential for hidden pattern detection in correlated data and efficiency in related optimization tasks.[4, 5, 6, 7, 8, 9, 10, 11, 12]
@@
Although we limit our scope to modeling and simulation in the following, it should be stressed that ML methods and ANNs have been widely employed in experimental studies (e.g., process regression, plasma diagnostics). For instance, as early as in 1992, they have been explored for parameter estimation in plasma etching based on optical emission spectroscopy (OES) and mass spectroscopy [13, 14] as well as using global process parameters paired with measured etch characteristics.[15] Plasma virtual metrology (VM) was conducted based on multivariate sensor data,[16] with regard to real-time fault detection in reactive ion etching,[17] batch process characterization in semiconductor fabrication,[18] and a deep learning VM framework based on OES data [19] and ’plasma information’ descriptors.[20] Further examples have devised an inverse reconstruction of intrinsic plasma properties such as the electron energy distribution function from OES diagnostics data [21] as well as an active learning guided scheme based on Fourier transform infrared spectroscopy data for parameter space exploration.[22]
@@
In contrast to these previous examples which rely on experimental and diagnostics data, the focus of this paper will be on data-driven aspects of plasma modeling and simulation. We consider how these methods can complement and/or replace theoretical approaches, including low and atmospheric pressure LTPs. Specifically, in section 2, we initially attempt to provide a comprehensive review of data-driven low-temperature plasma modeling. This will specifically cover plasma physics, plasma chemistry, plasma-surface interactions (PSIs), as well as process control and design aspects. In section 3, we further relate this to the closely linked disciplines of plasma fusion research, and to more general connections in data science and fluid dynamics research. In this context, an assessment and perspective of future developments and opportunities in LTP research will be devised. Overall, the main focus will be on DL methods, but also computational statistics methods like Gaussian process regression (GPR) will be outlined. Finally in section 4, concluding remarks with a perspective of the current state of the art in adjacent disciplines and the foreseen requirements, potential applications, and future developments of ML in plasma modeling and simulation in the next years 5–10 years will be drawn.
@@ 2 Review
A general perspective on the current state of data-driven LTP science – among many other aspects – has been laid out in The 2022 Plasma Roadmap[9]. Despite the ample momentum that data science and ML have recently gained in LTP, the potential of a community effort for systematic data harvesting and exploration to initiate a form of ’plasma informatics’ has been identified. Kambara et al.[12] have further detailed this paradigm with a comprehensive summary of science-based, data-driven developments. Whereas a wide range of plasma processing technologies have been addressed, a number of exemplary use cases of data-driven plasma modeling and simulation will be revisited later. Moreover, in an overarching effort with a broad scope of the field, data-driven plasma science has been reviewed.[10] Among other aspects, the fundamentals of data science and its application to space and astronomical plasmas, fusion plasmas, and LTPs have been elaborated. Owing to its broad review scope, LTP processing and its applications have only been generally assessed. In contrast, Bonzanini et al.[11] elaborate in more detail on the fundamentals of ML for LTPs. Therein, the authors not only provide a comprehensive introduction into the concepts and a collection of ML methods with relevant links to the literature, but also underpin their theoretical foundations with an instructive set of applied LTP case studies.
@@ 2.1 Plasma physics
LTPs for plasma processing entail a diverse zoo of physical and chemical phenomena with complex interactions. Modeling of such systems must, therefore, comprise at least the governing mechanisms and their intrinsic coupling. For example, consider the spatiotemporal dynamics of charged species coupled to the electromagnetic fields. Data-driven modeling of LTP discharge characteristics has followed three prevalent routes: (i) Data-driven modeling of a component of a physical model, based on data sets from numerical discharge simulations and finally coupled to the overall numerical simulation. (ii) Surrogate modeling that pursues to represent the complete discharge dynamics. (iii) Direct embedding of the physical discharge model equations to be solved into a physics-informed ML procedure.
@@
Contributions to this topic have prevalently considered separable submodels as part of superordinate model consisting of several components. Plasma chemistry presents a specific case detailed later in subsection 2.2. PSI is similarly addressed separately in subsection 2.3. A major aspect to be considered in the following concerns the description of electromagnetic fields, in particular in electrostatics description. In this case, the Poisson equation presents an elliptic partial differential equation (PDE) subject to time-dependent boundary conditions (BCs). This problem stands out because of its computational effort and difficulties in an efficient parallel solution in multiple spatial dimensions (as the information of the whole domain is required at every iteration). Early work by Zhang et al.[23] on 2D Monte Carlo (MC) simulations of a PN junction (conceptually related to methods applied in LTP) has considered reference data from solving the inhomogeneous Poisson equation subject to Dirichlet BCs using a finite-difference (FD) method. A multi-scale ANN was trained using the internal and boundary values of the electric potential as ground truth. Good agreement paired with a more than 10 times speedup has been reported. In an attempt to transfer a similar concept to LTPs, Trieschmann et al.[24] have implemented a 1D self-supervised learning scheme using an FD discretization to establish a physics-based loss, as for instance proposed by Dissanayake et al.[25] The data-driven solver was further assessed within a particle-in-cell (PIC) simulation of two collisionless counterpropagating electron streams. A relative error of the order of was reported for the electric potential. However, as the force acting on each particle requires the electric field, evaluation of the first derivative implied significant distortion of the spatiotemporal dynamics. This finding has been corroborated by Aguilar et al.[26] who have reported similar findings using fully connected ANNs (a.k.a. multilayer perceptrons (MLPs)[27]) and convolutional neural networks (CNNs)[28]. A 2D solution to Poisson’s equation coupled to a particle balance equation for a double headed streamer simulation (in drift-diffusion approximation for electrons and otherwise no transport) has been pursued by Cheng et al.[29] (closely related to the work Özbay et al.[30]) Therein, a comparison of different ANN architectures was presented, specifically a multi-scale network and a U-Net configuration.[31] Whereas a significant speedup has been consistently reported, the accuracy of the predictions was observed to agree only for short term prediction (ns). Note that here a simple rectangular domain was considered with the addition of a standardization scheme to extend to variable domain sizes. Further, the loss was composed of the inner domain and Dirichlet boundary values, paired with a physics-based loss representing the Laplacian. Zeng et al.[32] have followed a different paradigm using physics-informed neural networks (PINNs)[33, 34]. Therein, physical coordinates are input to an ANN and the underlying PDE is implemented as a physics-informed loss function. Generally, this loss function may include a set of possibly coupled PDEs to be solved within the domain, the BCs, and the initial conditions in case of transient phenomena.[33, 34] Automatic differentiation is exploited to obtain the respective local error gradients with respect to the input variables.[35] It resembles a paradigm closely linked to the backpropagation algorithm[27] in training of ANNs and enables a mesh-less representation of the PDE solution. In the referenced work, the inhomogeneous Poisson equation with rectangular and circular dielectric objects was solved in two and three dimensions, subject to time-dependent Dirichlet and/or Neumann BCs. Note that the authors have initially included time as a variable to consider time-dependent BCs (but later reverted to a time-invariant formulation), although Poisson’s equation essentially poses a static PDE problem. Hence, the error metric was composed of a domain loss representing the Laplacian and the loss corresponding to Dirichlet BCs. Whereas the application to dielectric barrier discharges (DBDs) was proposed as use case, an integration into a complete LTP model remains due. An approach dedicated to the requirements of complex LTP simulations has been put forward by Siffa et al.[36] based on a combination of U-Net,[31] DenseNet,[37] and transformer[38, 39] network architectures, called TransDenseUNet. On a 2D Cartesian mesh, the non-overlapping domain information was encoded into types (plasma, dielectric, BCs), corresponding physical values (dielectric constant, charge density, boundary potential/normal derivative), and spatial coordinates. A feature scaling of the physical equation was proposed and all input/output quantities were normalized prior training. The data set was composed of a few physical reference cases for intended applications and a substantial amount (10k to 50k) of randomly generated data samples. The training loss was combined with various contributions (including a gradient loss and a structure similarity index metric (SSIM)[40]) to increase the accuracy and the smoothness of the electric potential prediction (and the associated electric field). The proposed data-driven model was demonstrated to successfully generalize on arbitrary geometries, BCs, and charge distributions – including realistic cases encountered in LTP processing. Also here, an implementation within a complete plasma model remains due.
@@
With the goal of studying the complex interaction of electromagnetic waves with plasma, Desai et al.[41] have suggested a U-Net architecture to predict the 2D distribution of the root mean square electric field based on plasma density distributions. The ground truth data set was obtained from finite-difference time-domain (FDTD) simulations. Interestingly, while a simple regularized L2 loss was used, the quality of the prediction was assessed using the SSIM and the average percentage error. Whereas overall good agreement with reference simulations was achieved, an issue generally noticeable with many ANN surrogate models was observed. Specifically, the convergence toward a reliable prediction is subject to a spectral bias. Lower frequency features converge significantly faster than higher frequency features. Consequently, it may be computationally intractable to train until convergence is achieved for all frequencies.[42]
@@
An early example by Verma et al.[43] has considered data sets from 2D time-resolved fluid simulations of capacitively coupled radio-frequency discharges to establish surrogate models that capture reduced spatiotemporal modes of operation. In contrast to physical model based solutions, the authors have suggested their ANN surrogate model to be equation-free. The corresponding reduced order description and its latent space representation was suggested for parameter space exploration. This may otherwise be infeasible due to computational requirements. Another approach to capacitively coupled plasma surrogate modeling by the Hamaguchi lab has been devised based on 1D fluid and particle-in-cell with Monte Carlo collisions (PIC/MCC) simulations of an Ar discharge.[44] Instead of a grid-based approach for training and prediction (where the whole spatial domain is regressed on at once), MLPs were trained on phase-average plasma quantities sampled at varying positions within the discharge. The dynamics of an inductively coupled plasma with additional radio frequency substrate bias has been implemented based on HPEM (The Hybrid Plasma Equipment Model)[45] plasma simulations by Ko et al.[46] 2D axisymmetric simulations of an Ar etch plasma were performed with a fixed geometry (rectangular mesh) and varying process conditions (pressure, source power, bias power at 1 MHz and 13.56 MHz. The devised data set was further used to train a multi-encoder/decoder ANN, which encodes the process variables and geometry into a latent representation to finally decode the corresponding phase-averaged discharge quantities. Note that although the geometry was fixed in the presented study, the proposed ANN architecture with dedicated geometry inputs (sources, material and properties) may also be capable of taking varying geometrical features (e.g., chamber aspect ratio, wafer diameter) into account. Due to its low computational costs, the ML surrogate model allowed for a rapid exploration and optimization of a uniform and sufficient etch rate using the multi-objective particle swarm optimization (MOPSO) algorithm.
@@
PINNs were further used to directly consider the equations governing the physics of LTP discharges. Similar to the previously mentioned PINN solution to Poisson’s equation, corresponding PDEs may be implemented in the error metric during training.[33, 34] Automatic differentiation is again exploited to obtain desired partial derivatives with respect to the input variables.[35] Early work in LTP has been proposed by Kawaguchi et al.[50] who have considered a PINN solution to the stationary Boltzmann equation subject to ionization and electron attachment processes. An MLP has been trained to predict the electron velocity distribution function (EVDF) following an exponential ansatz. The solution was validated for two reference gas mixtures, whereas the dependence on a variation of the reduced electric field was used as a case study. Excellent agreement between the data-driven approach and MC simulations was reported. Several improvements to their initial scheme have been suggested by the authors. Specifically, a nondimensionalization of the equation combined with an improved ANN architecture was used.[51] The last is a fully connected ANN which includes two transformer layers for an improved training of the PINN.[52, 38] Further, a scheme to take into account variable reduced electric field values was proposed.[53] In both studies, overall good agreement with MC simulations was demonstrated. For the Ar gas 8 processes were included: momentum transfer, six electronic excitations and ionization. For the SF case 9 processes were included: momentum transfer, vibrational and electronic excitations, electron attachment and ionization. An extension to the time-dependent Boltzmann equation and assessment of the memory capacity of the ANN of 0.1% compared to mesh-based DNS calculations of EVDFs in 3D velocity space was further argued.
@@
In the frame of a PINN framework for LTP, Zhong et al.[54] have initially investigated stationary and transient 1D arc discharges. A fully connected ANN was used to implement the stationary Elenbaas–Heller equation (with decay term) as well as the time-dependent Elenbaas–Heller equation (considering mass and energy conservation). Whereas the PDE was generally solved using PINNs, the thermodynamic, transport, and radiation properties of the plasma were (partially) taken into account via so-called coefficient subnets. These resemble small regression networks to interpolate the corresponding plasma properties. In a follow-up study,[55] the authors have introduced this framework more generally in the frame of Coefficient-Subnet Physics-Informed Neural Network (CS-PINN) and Runge–Kutta Physics-Informed Neural Network (RK-PINN). For the first approach, a comparison with cubic spline interpolation was included. The last approach embeds PINNs in an implicit Runge–Kutta formalism. In a case study for three examples of Boltzmann equation (with ionization, electron attachment, elastic collision, and excitation), Poisson-drift-diffusion equations (for electrons and ions), and the previously addressed 1D arc discharge using the time-dependent Elenbaas–Heller equation, excellent agreement was noted. Zhong et al.[56] have lastly revisited the data-driven 1D arc discharge simulation, introducing the concept of meta-learning.[57] Therein, a meta loss was aggregated over multiple tasks (e.g., parameters governing discharge conditions, time-dependence) to enable an efficient two-step optimization and, therefore, reduce the training time. For each training task, a regular PINN was trained. In contrast, the meta network was trained to provide good initial weights for a PINN solving a new task. Based on a variation of arc discharge parameters and corresponding meta-training, a substantial acceleration with a factor in the range of 1.1 to 6.9 was reported.
@@
A different approach to PINNs following Raissi et al.[33] involves the discovery of PDEs from observations. A single PDE or a set of PDEs may be formulated as prototype equations. Therein, unknown coefficients determine the exact form of the equations. The problem can be phrased such that training a PINN not only yields the desired solution to the PDE, but includes the coefficients as parameters of the PINN. This concept has been used by Kawaguchi et al.[58] to “discover” electron transport coefficients from electron swarm maps measured by scanning drift-tube experiments. The temporal and spatial evolution of the electron swarm was formulated in two PDEs with corresponding transport coefficients. A fully connected ANN with transformer layers was utilized similar to previous works by the authors.[51] The tunable parameters are optimized along with ANN, without assuming any analytical form of the electron swarm map. It was concluded that more accurate transport coefficients could be obtained, including higher-order dependencies if required.

### v2

# Review: Machine learning for advancing low-temperature plasma modeling and simulation

@@ Abstract
Machine learning has had an enormous impact in many scientific disciplines. Also in the field of low-temperature plasma modeling and simulation it has attracted significant interest within the past years. Whereas its application should be carefully assessed in general, many aspects of plasma modeling and simulation have benefited substantially from recent developments within the field of machine learning and data-driven modeling. In this survey, we approach two main objectives: (a) We review the state-of-the-art, focusing on approaches to low-temperature plasma modeling and simulation. By dividing our survey into plasma physics, plasma chemistry, plasma-surface interactions, and plasma process control, we aim to extensively discuss relevant examples from literature. (b) We provide a perspective of potential advances to plasma science and technology. We specifically elaborate on advances possibly enabled by adaptation from other scientific disciplines. We argue that not only the known unknowns, but also unknown unknowns may be discovered due to the inherent propensity of data-driven methods to spotlight hidden patterns in data.
@@ 1 Introduction
Low-temperature plasmas (LTPs) consist of a mixture of neutral and charged species in thermal non-equilibrium. The governing discharge physics and chemistry lead to a zoo of (excited) species interacting with bounding surfaces. Their modification (e.g., activation, etching, functionalization) is exploited in the frame of plasma processing (e.g., plasma-enhanced etching, deposition, catalysis). LTPs also render a dominant part of the many steps required in semiconductor device manufacturing. This may involve a diverse portfolio of plasma discharges, chemical precursors, surface materials, and process recipe steps to achieve desired deposit/etch patterns.[1, 2] It may as well include indirect plasma discharges for example in the generation of extreme ultraviolet light for state of the art 13.5 nm wavelength lithography, required for device feature scales of the order of 10 nm and beyond (for instance termed 3 nm nodes).[3] Modeling and numerical simulation contribute an integral component in the design and optimization of related plasma processes. In the recent past decade, data-driven methods such as machine learning (ML), deep learning (DL), and artificial neural networks (ANNs), among other computational statistics methods, have (re-)gained enormous interest – which may be subsumed under the umbrella term artificial intelligence (AI). Although these fields have seen a renaissance and have continued to rapidly develop in the past decade, many fundamental concepts have been well established in computational statistics and related disciplines. The past and present interest in the context of plasma processing may be attributed to the potential for hidden pattern detection in correlated data and, particularly, the efficiency in related optimization tasks.[4, 5, 6, 7, 8, 9, 10, 11, 12]
@@
Although we limit our scope to modeling and simulation in the following, it should be stressed that ML methods and ANNs have been widely employed in experimental studies (e.g., process regression, plasma diagnostics). For instance, as early as in 1992, they have been explored for parameter estimation in plasma etching based on optical emission spectroscopy (OES) and mass spectroscopy, [13, 14] or global process parameters paired with measured etch characteristics.[15] Plasma virtual metrology (VM) was conducted based on multivariate sensor data,[16] with regard to real-time fault detection in reactive ion etching,[17] batch process characterization in semiconductor fabrication,[18] and a deep learning VM framework based on OES data [19] and ’plasma information’ descriptors.[20] Further examples have devised an inverse reconstruction of intrinsic plasma properties such as the electron energy distribution function from OES diagnostics data [21] as well as an active learning guided scheme based on Fourier transform infrared spectroscopy data for parameter space exploration.[22]
@@
In contrast to these previous examples which rely on experimental and diagnostics data, the focus of this paper will be on data-driven aspects of plasma modeling and simulation. We consider how these methods can complement and/or replace theoretical approaches, including low and atmospheric pressure LTPs. Specifically, in section 2, we initially attempt to provide a state-of-the-art review of data-driven LTP modeling. This specifically covers plasma physics, plasma chemistry, plasma-surface interactions (PSIs), as well as process control and design aspects. In section 3, we further relate this to the closely linked disciplines of plasma fusion research, and to more general connections in data science and fluid dynamics research. In this context, an assessment and perspective of future developments and opportunities in LTP research is devised. Overall, the main focus is on DL methods, but also computational statistics methods like Gaussian process regression (GPR) are outlined. Finally in section 4, concluding remarks with a perspective of the current state of the art in adjacent disciplines and the foreseen requirements, potential applications, and future developments of ML in plasma modeling and simulation in the next years 5–10 years is drawn.
@@ 2 Review
A general perspective on the current state of data-driven LTP science – among many other aspects – has been laid out in The 2022 Plasma Roadmap[9]. Despite the ample momentum that data science and ML have recently gained in LTP, the potential of a community effort for systematic data harvesting and exploration to initiate a form of ’plasma informatics’ has been identified therein. Kambara et al.[12] have further detailed this paradigm with a comprehensive summary of science-based, data-driven developments. A wide range of plasma processing technologies have been addressed, whereas a number of exemplary use cases of data-driven plasma modeling and simulation will be revisited later. Moreover, in an overarching effort the broad scope of the field of data-driven plasma science has been reviewed.[10] Among other aspects, the fundamentals of data science and its application to space and astronomical plasmas, fusion plasmas, and LTPs have been elaborated. Owing to its broad review scope, LTP processing and its applications have only been generally assessed. In contrast, Bonzanini et al.[11] elaborate in more detail on the fundamentals of ML for LTPs. Therein, the authors not only provide a comprehensive introduction into the concepts and a collection of ML methods with relevant links to the literature, but also underpin their theoretical foundations with an instructive set of applied LTP case studies.
@@ 2.1 Plasma physics
LTPs for plasma processing entail a diverse zoo of physical and chemical phenomena with complex interactions. Modeling of such systems must, therefore, comprise at least the governing mechanisms and their intrinsic coupling. For example, consider the spatiotemporal dynamics of charged species coupled to electromagnetic fields. Data-driven modeling of LTP discharge characteristics has followed three prevalent routes: (i) Data-driven modeling of a component of a physical model, based on data sets from numerical discharge simulations and finally coupled to the overall numerical simulation. (ii) Surrogate modeling that pursues to represent the complete discharge dynamics. (iii) Direct embedding of the physical discharge model equations to be solved into a physics-informed ML procedure.
@@
Contributions to this topic have prevalently considered separable submodels as part of a superordinate model consisting of several components. Plasma chemistry presents a specific case detailed later in subsection 2.2. PSI is similarly addressed separately in subsection 2.3. A major aspect to be considered in the following concerns the description of electromagnetic fields, in particular in electrostatics description. In this case, the Poisson equation presents an elliptic partial differential equation (PDE) subject to time-dependent boundary conditions (BCs). This problem stands out because of its computational effort and difficulties in an efficient parallel solution in multiple spatial dimensions (as the information of the whole domain is required at every iteration). Early work by Zhang et al.[23] on 2D Monte Carlo (MC) simulations of a PN junction (conceptually related to methods applied in LTP) has considered reference data from the solution of the inhomogeneous Poisson equation subject to Dirichlet BCs using a finite-difference (FD) method. A multi-scale ANN was trained using the internal and boundary values of the electric potential as ground truth. Good agreement paired with a more than 10 times speedup has been reported. In an attempt to transfer a similar concept to LTPs, Trieschmann et al.[24] have implemented a 1D self-supervised learning scheme using an FD discretization to establish a physics-based loss, as for instance proposed by Dissanayake et al.[25] The data-driven solver was further assessed within a particle-in-cell (PIC) simulation of two collisionless counterpropagating electron streams. A relative error of the order of was reported for the electric potential. However, as the force acting on each particle requires the electric field, evaluation of the first derivative implied significant distortion of the long-term spatiotemporal dynamics. This finding has been corroborated by Aguilar et al.[26] who have reported similar findings using fully connected ANNs (a.k.a. multilayer perceptrons (MLPs)[27]) and convolutional neural networks (CNNs)[28]. A 2D solution to Poisson’s equation coupled to a particle balance equation for a double headed streamer simulation (in drift-diffusion approximation for electrons and otherwise no transport) has been pursued by Cheng et al.[29] (closely related to the work Özbay et al.[30]) Therein, a comparison of different ANN architectures was presented, specifically a multi-scale network and a U-Net configuration.[31] Whereas a significant speedup has been consistently reported, the accuracy of the predictions was observed to agree only for short term prediction ( ns). Note that here a simple rectangular domain was considered with the addition of a standardization scheme to extend to variable domain sizes. Further, the loss was composed of the inner domain and Dirichlet boundary values, paired with a physics-based loss representing the Laplacian. Zeng et al.[32] have followed a different paradigm using physics-informed neural networks (PINNs)[33, 34]. Therein, physical coordinates are input to an ANN and the underlying PDE is implemented as a physics-informed loss function. Generally, this loss function may include a set of possibly coupled PDEs to be solved within the domain, the BCs, and the initial conditions in case of transient phenomena.[33, 34] Automatic differentiation is exploited to obtain the respective local error gradients with respect to the input variables.[35] It resembles a paradigm closely linked to the backpropagation algorithm[27] for training ANNs and enables a mesh-less representation of the PDE solution. In the referenced work, the inhomogeneous Poisson equation with rectangular and circular dielectric objects was solved in two and three dimensions, subject to time-dependent Dirichlet and/or Neumann BCs. Note that the authors have initially included time as a variable to consider time-dependent BCs (but later reverted to a time-invariant formulation), although Poisson’s equation essentially poses a static PDE problem. Hence, the error metric was composed of a domain loss representing the Laplacian and the loss corresponding to Dirichlet BCs. Whereas the application to dielectric barrier discharges (DBDs) was proposed as use case, an integration into a complete LTP model remains due. An approach dedicated to the requirements of complex LTP simulations has been put forward by Siffa et al.[36] based on a combination of U-Net,[31] DenseNet,[37] and transformer[38, 39] network architectures, called TransDenseUNet. On a 2D Cartesian mesh, the non-overlapping domain information was encoded into types (plasma, dielectric, BCs), corresponding physical values (dielectric constant, charge density, boundary potential/normal derivative), and spatial coordinates. A feature scaling of the physical equation was proposed and all input/output quantities were normalized prior training. The data set was composed of a few physical reference cases for intended applications and a substantial amount (10k to 50k) of randomly generated data samples. The training loss was combined with various contributions (including a gradient loss and a structural similarity index measure (SSIM)[40]) to increase the accuracy and the smoothness of the electric potential prediction (and the associated electric field). The proposed data-driven model was demonstrated to successfully generalize on arbitrary geometries, BCs, and charge distributions – including realistic cases encountered in LTP processing. Also here, an implementation within a complete plasma model remains due.
@@
With the goal of studying the complex interaction of electromagnetic waves with plasma, Desai et al.[41] have suggested a U-Net architecture to predict the 2D distribution of the root mean square electric field based on plasma density distributions. The ground truth data set was obtained from finite-difference time-domain (FDTD) simulations. Interestingly, while a simple regularized L2 loss was used, the quality of the prediction was assessed using the SSIM and the average percentage error. Whereas overall good agreement with reference simulations was achieved, an issue generally noticeable with many ANN surrogate models was observed. Specifically, the convergence toward a reliable prediction is subject to a spectral bias. Lower frequency features converge significantly faster than higher frequency features. Consequently, it may be computationally intractable to train until convergence is achieved for all relevant frequencies.[42]
@@
An early example by Verma et al.[43] has considered data sets from 1D time-resolved fluid simulations of capacitively coupled radio-frequency discharges to establish surrogate models for integrated quantities such as the plasma sheath voltages/currents, that capture reduced temporal modes of operation. In contrast to physical model based solutions, the authors have highlighted the equation-free representation of their long short-term memory (LSTM, cf. subsection 3.1)) surrogate model. The corresponding reduced order description and its latent space representation was suggested for parameter space exploration. This may otherwise be infeasible due to computational requirements. Another approach to capacitively coupled plasma surrogate modeling by Hamaguchi and co-workers has been devised based on 1D fluid and particle-in-cell with Monte Carlo collisions (PIC/MCC) simulations of an Ar discharge.[44] Instead of a grid-based approach for training and prediction (where the whole spatial domain is regressed on at once), MLPs were trained on phase-average plasma quantities sampled at varying positions within the discharge. The dynamics of an inductively coupled plasma with additional radio frequency substrate bias has been implemented based on HPEM (hybrid plasma equipment model)[45] simulations by Ko et al.[46] 2D axisymmetric simulations of an Ar etch plasma were performed with a fixed geometry (rectangular mesh) and varying process conditions (pressure, source power, bias power at 1 MHz and 13.56 MHz). The devised data set was further used to train a multi-encoder/decoder ANN, which encodes the process variables and geometry into a latent representation to finally decode the corresponding phase-averaged discharge quantities. Note that although the geometry was fixed in the presented study, the proposed ANN architecture with dedicated geometry inputs (sources, material and properties) may also be capable of taking varying geometrical features (e.g., chamber aspect ratio, wafer diameter) into account. Due to its low computational costs, the ML surrogate model allowed for a rapid exploration and optimization of a uniform and sufficient etch rate using the multi-objective particle swarm optimization algorithm.
@@
PINNs were further used to directly consider the equations governing the physics of LTP discharges. Similar to the previously mentioned PINN solution to Poisson’s equation, corresponding PDEs may be implemented in the error metric during training.[33, 34] Automatic differentiation is again exploited to obtain desired partial derivatives with respect to the input variables.[35] Early work in LTP has been proposed by Kawaguchi et al.[50] who have considered a PINN solution to the stationary Boltzmann equation subject to ionization and electron attachment processes. An MLP has been trained to predict the electron velocity distribution function (EVDF) following an exponential ansatz. The solution was validated for two reference gas mixtures, whereas the dependence on a variation of the reduced electric field was used as a case study. Excellent agreement between the data-driven approach and MC simulations was reported. Several improvements to their initial scheme have been suggested by the authors. Specifically, a nondimensionalization of the equation combined with an improved ANN architecture was used.[51] The last is a fully connected ANN which includes two transformer layers for an improved training of the PINN.[52, 38] Further, a scheme was proposed to take into account variable reduced electric field values .[53] In both studies, overall good agreement with MC simulations was demonstrated. For Ar gas eight processes were included: momentum transfer, six electronic excitations and ionization. For SF gas nine processes were included: momentum transfer, vibrational and electronic excitations, electron attachment and ionization. An extension to the time-dependent Boltzmann equation and assessment of the memory capacity of the ANN of 0.1% compared to mesh-based DNS calculations of EVDFs in 3D velocity space was further argued.
@@
In the frame of a PINNs for LTP, Zhong et al.[54] have initially investigated stationary and transient 1D arc discharges. A fully connected ANN was used to implement the stationary Elenbaas–Heller equation (with decay term) as well as the time-dependent Elenbaas–Heller equation (considering mass and energy conservation). Whereas the PDE was generally solved using PINNs, the thermodynamic, transport, and radiation properties of the plasma were (partially) taken into account via so-called coefficient subnets. These resemble small regression networks to interpolate the corresponding plasma properties. In a follow-up study,[55] the authors have introduced this framework more generally in the frame of Coefficient-Subnet Physics-Informed Neural Network (CS-PINN) and Runge–Kutta Physics-Informed Neural Network (RK-PINN). For the first approach, a comparison with cubic spline interpolation was included. The last approach embeds PINNs in an implicit Runge–Kutta formalism. In a case study for three examples of Boltzmann equation (with ionization, electron attachment, elastic collision, and excitation), Poisson-drift-diffusion equations (for electrons and ions), and the previously addressed 1D arc discharge using the time-dependent Elenbaas–Heller equation, excellent agreement was noted. Zhong et al.[56] have lastly revisited the data-driven 1D arc discharge simulation, introducing the concept of meta-learning.[57] Therein, a meta loss was aggregated over multiple tasks (e.g., parameters governing discharge conditions, time-dependence) to enable an efficient two-step optimization and, therefore, reduce the training time. For each training task, a regular PINN was trained. In contrast, the meta network was trained to provide good initial weights for a PINN solving a new task. Based on a variation of arc discharge parameters and corresponding meta-training, a substantial acceleration with a factor in the range of 1.1 to 6.9 was reported.
@@
A different approach to PINNs following Raissi et al.[33] involves the discovery of PDEs from observations. A single PDE or a set of PDEs may be formulated as prototype equations. Therein, unknown coefficients determine the exact form of the equations. The problem can be phrased such that training a PINN not only yields the desired solution to the PDE, but includes the coefficients as parameters of the PINN. This concept has been used by Kawaguchi et al.[58] to “discover” electron transport coefficients from electron swarm maps measured by scanning drift-tube experiments. The temporal and spatial evolution of the electron swarm was formulated in two PDEs with corresponding transport coefficients. A fully connected ANN with transformer layers was utilized similar to previous works by the authors.[51] The tunable parameters are optimized along with the ANN, without assuming any analytical form of the electron swarm map. It was concluded that more accurate transport coefficients could be obtained, including higher-order dependencies if required.